POTENTIAL ENERGY SURFACES AND VIBRATIONAL ENERGY LEVELS OF DCCl AND HCCl IN THEIR THREE LOW-LYING STATES^{a}

HUA-GEN YU, TREVOR J. SEARS and \underline{JAMES T. MUCKERMAN}, Chemistry Department, Brookhaven National Laboratory, Upton, NY 11973-5000 USA.

We present ab initio multi-reference configuration interaction (MRCI) calculations of potential energy surfaces of HCCl in its three low-lying electronic states (X̃^{1}A′, ã^{3}A′′ and Ã^{1}A′′) and for the spin-orbit coupling between the X̃ and ã states. The two singlet states become a degenerate^{1}Δ state in collinear geometries. The potential energy surfaces are interpolated from 6075 MRCI energy points. The final surfaces are slightly adjusted using a coordinate and energy scaling approach. The T_{e} values of the ã^{3}A′′ and Ã^{1}A′′ states are computed to be 2122.0 and 12209.8 cm^{−1}, respectively. Vibrational energy levels of the three states of DCCl and HCCl taking into account the Renner-Teller effect and spin-orbit coupling are computed. The calculated vibronic energy levels are in good agreement with the available experimental values.

aWork at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.